# pyCFS-data: Data Processing Framework in Python for openCFS

### v1

@@ 2.1 Reader class (CFSReader)
, which reads the file file.cfs into the data objects mesh and results, containing the description of the computational grid and result data, respectively.
@@ 2.2 Writer class (CFSWriter)
f.create_file(mesh_data=mesh, result_data=results)
@@
2.3 Data structures (CFSMeshData,CFSResultData, etc.)
@@
CFSResultData: Data structure containing result data (one object currently supports a single MultiStep only!). Its attributes and properties are depicted in Fig. 4.
Figure 4: Structure of CFSResultData class. Attributes are highlighted in orange, Properties in purple, and respective data types in grey.
@@
Field data can be defined on the nodes or the cell centroids of a computational grid (ResType attribute) in the time or the frequency domain (AnalysisType attribute).
@@
3 Operators for data processing (opertors)
@@ 3.1.1 Node2Cell
$e_{\square}=\sum_{i=1}^{\mathrm{n}}v_{i}\,.$ (1)
Thereby, $e_{\square}$ is the load located to the cell, $\mathrm{n}$ is the number of nodes of one element, and $v_{i}$ is the nodal loads. The following example shows this methodology by considering one linear triangular element:
Figure 7: Cell to node interpolator.
Parameters include
  • coordinates: Node coordinate array of source mesh.
  • connectivity: Element connectivity array of source mesh.
@@ 3.1.2 Cell2Node
$v_{\square}=\frac{1}{\mathrm{n}}e_{i}\,.$ (2)
@@
Figure 8: Node to cell interpolator.
Parameters include
  • coordinates: Node coordinate array of source mesh.
  • connectivity: Element connectivity array of source mesh.
@@ 3.1.3 Nearest Neighbour
Parameters include
  • source_coord: Coordinate array of source positions.
  • target_coord: Coordinate array of target positions.
  • num_neighbors: Number of considered Nodes.
  • interpolation_exp: Exponent for calculation of interpolation weight function.
  • formulation: Search direction for nearest neighbor search. By default, the direction is chosen automatically based on a number of source and target points.
@@ 3.2 Transformation operations (transformation)
3.2.1 Fit geometry
@@
Parameters include
  • mesh_data_fit: CFSMeshData object of grid to fit
  • result_data_fit: CFSResultData object of vector data to transform with mesh fit
  • mesh_data_target: CFSMeshData object of target grid
  • transform_param_init: Initial transformation parameters
@@ 4.1 Ansys Mechanical (ansys_io)
  • Reading Ansys result files (*.rst) into CFSMeshData and CFSResultData objects.
@@ 4.2 EnSight Case Gold (ensight_io)
Module containing data processing utilities for reading EnSight Case Gold files. Highlights currently include:
@@
  • Conversion to CFSMeshData and CFSResultData objects.
@@ 4.3 Polytec PSV export data (psv_io)
  • Conversion to and from CFSMeshData and CFSResultData objects into sdypy-EMA compatible dict format.
@@ 4.4 NiHu Matlab export data (nihu_io)
Module containing data processing utilities for reading NiHu structures as Matlab export files. Highlights currently include:

### v2

@@ 2 I/O of openCFS based HDF5 files (io)
Module for input and output operations of openCFS type mesh and result data files. The simplest way of reading and writing files is via the top level utility functions read_file, read_mesh, read_data, write_file. A simple example reads
from pyCFS.data import io
mesh, results = io.read_file(’file.cfs’)
io.write_file(’file_out.cfs’, mesh, results)
, reading a file and writing the content to a new file. Figure 1 gives an overview of the I/O module’s structure.
Figure 1: Organization of classes for data structures.
@@ 2.1 Reader class (CFSReader)
, which reads the file file.cfs into the data objects mesh and results, containing the description of the computational grid and all result data of the default analysis step (default is $1$), respectively.
@@ 2.2 Writer class (CFSWriter)
f.create_file(mesh=mesh, result=results)
@@
2.3 Data structures (CFSMeshData,CFSResultContainer, etc.)
@@
CFSResultContainer: Data structure containing result data (one object currently supports a single MultiStep only!). Its attributes and properties are depicted in Fig. 4.
Figure 4: Structure of CFSResultContainer class. Attributes are highlighted in orange, Properties in purple, and respective data types in grey.
@@
Field data can be defined on the nodes or the cell centroids of a computational grid, while history data can also be defined on Region (ResType attribute), defining the required shape of the array data. Field data requires $(N,M,D)$ while history data requires $(N,D)$, with the number of steps $N$, the number of degrees of freedom (DOF) $M$ and the number of dimensions $D$. Data can be defined in the time or the frequency domain, based on the AnalysisType attribute.
[…]
Dimension names can be defined optionally (DimNames attribute). If not specified, they are chosen automatically based on the shape of the data array.
[…]
The attribute IsComplex can be defined optionally to ensure writing complex-valued data or enforce real-valued data. If not specified, it will be chosen automatically based on the corresponding analysis type (AnalysisType attribute).
@@
3 Operators for data processing (operators)
@@ 3.1.1 Node2Cell
$e_{\square}=\frac{1}{n}\sum_{i=1}^{n}v_{i}\,.$ (1)
Thereby, $e_{\square}$ is the data assigned to the element, $\mathrm{n}$ is the number of nodes of the element, and $v_{i}$ are the nodal loads.
Figure 7: Node to cell interpolator.
@@ 3.1.2 Cell2Node
$v_{\square}=\frac{1}{n}\sum_{i=1}^{n}e_{i}\,.$ (2)
Thereby, $v_{\square}$ is the data located to the node, $n$ the number of adjacent elements to the node, and $e_{i}$ the data of cell $i$.
@@
Figure 8: Cell to node interpolator.
@@ 3.1.3 Nearest Neighbour
Two search directions for the nearest neighbor search are implemented:
@@ 3.1.4 Projection-based interpolation
3.2 Modal analysis tools (modal_analysis)
Common metrics used in modal analysis including the Modal Assurance Criterion (MAC), the Modal Scale Factor (MSF), and the Modal Complexity Factor (MCF).
3.3 Synthetic Noise Generation and Radiation (SNGR) source computation (sngr)
Computation of synthetic flow acoustic sources based on the theory of C. Bailly [1] for the SNGR similar to the implementation described in [5].
@@ 3.4 Transformation operations (transformation)
3.4.1 Extrude/Revolve mesh(extrude_mesh_region, revolve_mesh_region)
Create a 3D/2D mesh by extruding/revolving a 2D/1D base mesh along a path.
3.4.2 Transform mesh data (transform_mesh_data)
Perform a rigid body transformation of a mesh and its corresponding vector data.
3.4.3 Fit geometry (fit_mesh)
@@ 4.1 Ansys Mechanical (ansys_io)
  • Reading Ansys result files (*.rst) into CFSMeshData and CFSResultContainer objects.
@@
4.2 CGNS reader (cgns_io)
Module for reading mesh files in CGNS file format, e.g. exported from ANSYS Mechanical.
[…]
Only HDF5 based CGNS files can be read. Older CGNS versions also utilise the ADF file format, which must be converted to HDF5 beforehand. This can be done e.g. on a Linux system using cgnslib with the command: adf2hdf \$CGNS_FILE
@@ 4.3 EnSight Case Gold (ensight_io)
Module containing data processing utilities for reading EnSight Case Gold files, e.g. exported from common CFD software like openFOAM, Ansys Fluent, or StarCCM+. Highlights currently include:
@@
  • Conversion to CFSMeshData and CFSResultContainer objects.
[…]
This submodule make use of VTK, which is not a standard requirement of pyCFS. The additional dependencies can be installed via pip using pip install pycfs[vtk].
4.4 Exodus file format (exodus_io)
Module for reading mesh files in Exodus file format, e.g. generated with Coreform Cubit.
@@ 4.5 Polytec PSV export data (psv_io)
  • Conversion to and from CFSMeshData and CFSResultContainer objects into sdypy-EMA compatible dict format.
@@ 4.6 NiHu Matlab export data (nihu_io)
Module containing data processing utilities for reading NiHu [2] Matlab structures. Highlights currently include:
@@
4.7 STL reader (stl_io)
Function for reading mesh data from STL files.